OpenStudy (anonymous):

A biologist who studies mice is looking at how many offspring they produce in a month. At the end of the study, 27 mice have 0-5 babies, 23 have 6-10 babies, 32 mice have 11-15 babies, and 18 have 16-20 babies. Construct an experimental probability table and represent your probabilities in percentage form. What is the experimental probability that a mouse will have over 10 babies

OpenStudy (kropot72):

Total number of mice in the study = 27 + 23 + 32 + 18 = 100  Can you complete the table?

OpenStudy (anonymous):

Ihave no idea how to, this is a question to see how much i know about the class before i take it so i was seeing how i would do this

OpenStudy (anonymous):

I wanna learn how to if you can help me?

OpenStudy (anonymous):

6-10 would be 23 then 11-15=32 and 16-20=18 then the tables done? & thats all that there is to the question to be answered

OpenStudy (kropot72):

 Now you need to find the percentage that have more than 10 babies. So just add the percentages in the two right hand columns of the table.

OpenStudy (anonymous):

50 ?

OpenStudy (kropot72):

Yes, 50 % is correct.

OpenStudy (anonymous):

Thank you that was really easy, when i see word problems i get confused lol i hate them!

OpenStudy (anonymous):

How could i do this? Create your own probability problem using a number cube. Classify your event and its compliment. Are the events mutually exclusive? Find the probability of the event occurring, its complement occurring, the event and its compliment occurring, and the event or its complement occurring.